# CORT: Class-Oriented Real-time Tracking for Embedded Systems

### v1

@@ IV-B Comparison with BYTEtrack
In this case, because of the progressive overlap, the position of the two objects becomes rapidly imprecise due to the absence of a measure to correct the Kalman filter prediction. During this transient, the bus can still be matched, while the the two cars under occlusion are updated by dead reckoning. When the occluded objects becomes visible again, they reappear in a position that is far from the last detection. Since BYTEtrack only uses the IoU match, the chance of missing the recovery of the track is high, especially for small objects. In fact, if the motion of the hidden objects results to be highly non-linear during the occlusion, its predicted position is likely to have a negligible IoU with the detected one or a high IoU with the wrong track, possibly causing an id-switch. In contrast, the solution implemented in CORT can recover the track in the second stage of the cascade matching, because it exploits the features extracted by the re-identification network. Then, using the cosine similarity, it is able to get a more reliable association score, so recognizing a previously occluded object, rather than creating a new track. The situation described above is very common in multi-object tracking and the results reported in Table II confirm that the proposed solution improves the overall performance with respect to BYTEtrack by significantly reducing the number of id switches, with a small increase in the inference time.
@@
Proposed solution 17.8 3004 72.30 79.126
BYTEtrack 19.1 22989 62.79 80.329
@@
TABLE II: A comparison of the BYTEtrack and the proposed class-based solution using different evaluation metrics.
@@
For the sake of fairness, the results reported in Table II have been obtained by replacing the detection network of BYTEtrack with the one used in the proposed architecture, which is a YOLOv8 trained on the COCO dataset [42]. In this way, both trackers worked with the same set of detections. The achieved results also show that the re-identification stage slightly increases the execution time of the proposed tracker with respect to BYTEtrack. To evaluate such a difference, however, one should consider that BYTEtrack is optimized for tracking multiple objects belonging to a single class, while the proposed model is optimized for multi-classes tracking.
IV-C Behavior under unstable detections
@@
This section examines the behavior of the standard and the multi-class approach in two critical situations in which the detection network could provide sporadic erroneous outputs in a frame sequence. To show the potential impact of such events on the performance of the tracker, consider a scenario similar to the one illustrated in Figure 7, where the red car appears in the scene at time $t_{a}$, but it is misclassified as a bus at time $t_{a}+2$. Also assume that all the other objects are correctly classified across the sequence. For the sake of clarity, only the detection responsible for the problem is shown in all the frames.
@@
To better understand the negative impact of such misclassification, Figure 7 also reports the evolution of the Hungarian matrices for both the standard monolithic approach and the proposed class-based solution, highlighting for each frame how this specific detection affects the size of the matrices and, hence, the number of operations required to fill them.
Note that, at time $t_{a}$, there are four predictions and five detections, due to the arrival of the new entity. As described in Section III-A, this requires the creation of a dummy row in the cost matrix to start the initialization phase of the new entity. Four of these five detections will be matched by the Hungarian algorithm with the predictions related to the pre-existing objects. The remaining one is used to create the data structure for the new track. In all trackers, the initialization phase requires a given amount of subsequent frames in which the new object must match with a detection. Such a transient is used to reject sporadic false detections and is also needed to initialize the covariance matrix of the Kalman Filter and adapt it to the motion of the new entity. If a new object is misclassified in such a transient, the track under initialization does not satisfy the requirements to be promoted as an active track. The result is that the object will be tracked with a larger delay, or its tracklet will never be activated, if the problem continues to occur. Note that, although this problem is common to all tracking algorithms, the multi-class solution is less affected in terms of computation with respect to the monolithic approach. In fact, in the multi-class match, a class change increases the dimension of two small matrices by one, whereas in the monolithic approach, the dimension of the Hungarian matrix is always increased by two.
Consider now a slightly different scenario in which all the objects in the scene are correctly tracked. Suppose that, for a given frame, the following two events occur for two close objects at the same time:
  • Missed detection: one of the object is not detected in the frame.
  • Wrong classification: the detection having an IoU value greater than zero with the two predictions changes its class.
Under this assumption, we have two predictions of different classes, but only one detection that can be used for updating the Kalman filter. If the detection contended between the two predictions changes its class faster than the number of frames tolerated by the tracker to remove a track, then a track that should have been eliminated can persist longer than needed. Such a persistent track also causes additional operations to be performed that could have been avoided. Notice that, also in this scenario, the proposed class-based solution behaves better than the standard monolithic approach. In fact, since the standard method only uses one matrix, maintaining a track active is more expensive in terms of operations, and this becomes worse as the number of objects in the scene increases. Vice versa, the multi-class approach requires fewer computations because the class matrices are always smaller than the one in the monolithic solution.
@@
A simple solution to attenuate the two problems described in this section is to increase the acceptance threshold of the detection network. Raising the value of such a threshold ensures that only reliable detections are passed to the tracker to compute the costs for the assignment problem. At the same time, raising the threshold increases the number of false negatives. However, this is not a major issue, because the Kalman filter can efficiently compensate for some missing detections. This is the reason why in Table II the value for the precision (MOTP) is lower than that of the BYTEtrack, while the accuracy (MOTA) is higher because of the reduced number of id-switches.
V Conclusions
@@
This work presented CORT, an efficient method to enhance the timing performance of modern real-time tracking algorithms. Such an enhancement is obtained by splitting the Hungarian matching algorithm into smaller association problems, carried out per class. Considering that the matching algorithm has an $O(n^{3})$ time complexity, where $n$ is the number of elements in the cost matrix, and that each sub-matrix can be solved in parallel on different cores, the proposed approach significantly improves the performance of state of the art tracking by detection systems. As highlighted in Section IV, as the number of objects to be tracked increases, the proposed parallel solution becomes more and more efficient than the monolithic one. This is also confirmed by the experimental results reported in Table I, showing that the gap between the execution time of the two methods significantly increases for tracking scenarios with a larger number of objects, highlighting the better scalability of the proposed solution.
As a future work, we plan to accelerate on FPGA some vectorizable operations of the proposed pipeline, as the computation of the association scores. Another optimization that can be further investigated is the dynamic tuning of the hyperparameters of the algorithm. In fact, as reported in Section IV-C, changing these values has a direct impact on the behavior of the tracker. A runtime adaptation would be desirable to maximize the amount of information extracted from each frame while avoiding the problems reported in Section IV.
VI Acknowledgments
@@
This work was partially supported by the Department of Excellence of Robotics and AI of the Sant’Anna School of Advanced Studies.

### v2

@@ 4.2 Comparison with BYTEtrack
In this case, because of the progressive overlap, the position of the two objects becomes rapidly imprecise due to the absence of a measure to correct the Kalman filter prediction. During this transient, the bus can still be matched, while the the two cars under occlusion are updated by dead reckoning. When the occluded objects becomes visible again, they reappear in a position that is far from the last detection. Since BYTEtrack only uses the IoU match, the chance of missing the recovery of the track is high, especially for small objects. In fact, if the motion of the hidden objects results to be highly non-linear during the occlusion, its predicted position is likely to have a negligible IoU with the detected one or a high IoU with the wrong track, possibly causing an id-switch. In contrast, the solution implemented in CORT can recover the track in the second stage of the cascade matching, because it exploits the features extracted by the re-identification network. Then, using the cosine similarity, it is able to get a more reliable association score, so recognizing a previously occluded object, rather than creating a new track. The situation described above is very common in multi-object tracking and the results reported in Table 2 confirm that the proposed solution improves the overall performance with respect to BYTEtrack by significantly reducing the number of id switches, with a small increase in the inference time.
@@
Proposed solution 23.8 3004 72.30 79.13
BYTEtrack 19.1 22989 62.79 80.33
@@
Table 2: A comparison of the BYTEtrack and the proposed class-based solution using different evaluation metrics.
@@
For the sake of fairness, the results reported in Table 2 have been obtained by replacing the detection network of BYTEtrack with the one used in the proposed architecture, which is a YOLOv8 trained on the COCO dataset [45]. In this way, both trackers worked with the same set of detections. The achieved results also show that the re-identification stage slightly increases the execution time of the proposed tracker with respect to BYTEtrack. To evaluate such a difference, however, one should consider that BYTEtrack is optimized for tracking multiple objects belonging to a single class, while the proposed model is optimized for multi-classes tracking. MOTA and MOTP were selected as evaluation metrics because they directly capture CORT’s goals of identity consistency and precise localization in multi-class tracking. MOTA consolidates tracking errors, like ID switches and occlusions, providing a straightforward measure of identity handling, while MOTP evaluates the alignment of tracked positions with the ground truth. HOTA, instead, is designed to capture long-term re-identification relations, which are not prioritized in the proposed solution.
4.3 Evaluation on KITTI 2D
The aim of this experiment is to show the results obtained by CORT on a standard multi-class multi-object tracking benchmark, like KITTI [46]. The KITTI dataset includes seven different classes, but the benchmark returns the score evaluation on only two of them: car and person. This choice was made by the benchmark creators because these two classes were considered to have a sufficient number of occurrences to ensure a reliable evaluation. Table 3 reports the scores obtained for the two classes by the proposed solution (CORT) and DeepSORT.
Results for class Car
Tracker IDS MOTA MOTP HOTA FPS CORT (two stages) 184 $\uparrow$ 85.66 $\uparrow$ 86.99 $\uparrow$ 72.64 $\uparrow$ 27.9 CORT (IoU only) 256 $\downarrow$ 73.51 $\downarrow$ 82.51 $\uparrow$ 70.31 $\uparrow$ 30 DeepSORT 237 78.14 76.66 59.98 19.7 BYTEtrack 242 76.11 78.37 71.63 24.8
Results for class Person
Tracker IDS MOTA MOTP HOTA FPS CORT (two stages) 263 $\uparrow$ 50.94 $\uparrow$ 66.45 $\uparrow$ 43.98 $\uparrow$ 27.2 CORT (IoU only) 348 $\downarrow$ 43.56 $\downarrow$ 57.85 $\downarrow$ 38.96 $\uparrow$ 30 DeepSORT 312 45.90 63.63 41.10 18.3 BYTEtrack 324 44.82 62.43 41.62 24.2
Table 3: Evaluation results on the Kitti 2D multi-object tracking benchmark.
The proposed solution demonstrates significant improvements in both MOTA and MOTP metrics compared to DeepSORT, achieving higher tracking accuracy and localization precision. For the Car class, the proposed solution achieves a MOTA of 85.66 and MOTP of 86.99, outperforming DeepSORT’s respective scores of 78.14 and 76.66. Similarly, for the Person class, the proposed method achieves a MOTA of 50.94 and MOTP of 66.45, confirming better results with respect to DeepSORT’s scores of 45.90 and 63.63.
CORT also demonstrates a balanced performance in terms of HOTA, outperforming DeepSORT and showing strengths in maintaining tracking consistency for dynamic objects. In particular, note that the HOTA improvement achieved by CORT over DeepSORT is higher for the Car class, confirming the effectiveness of the proposed solution in tracking objects with faster dynamics, whereas people are easier to re-associate after occlusions for both models.
We also carried out an ablation study to evaluate the effect of the second stage (in charge of re-identification) on the tracking speed and performance of CORT. The results reported in Table 3 remark the importance of the second stage for both reducing the number of ID switches (IDS) and recovering from temporary occlusions. Specifically, for the Car class, the IDS is reduced from 256 to 184, while MOTA improves from 73.51 to 85.66, and HOTA from 70.31 to 72.64. Similarly, for the Person class, the IDS is reduced from 348 to 263, with corresponding improvements in MOTA (from 43.56 to 50.94) and HOTA (from 38.96 to 43.98). These improvements come at the cost of a slight reduction in the frame rate, which decreases from 30 fps to 27.9 fps for the Car class and from 30 fps to 27.2 fps for the Person class. This feature provides an additional flexibility to the user for configuring the CORT pipeline depending on the requirements of the specific application. A direct comparison with BYTEtrack highlighted CORT’s advantages in both accuracy and efficiency. For the Car class, CORT achieved higher HOTA (72.64 vs. 71.63) and MOTA (85.66 vs. 76.11), with fewer ID switches (184 vs. 242) and a slightly higher FPS (27.9 vs. 24.8). Similarly, for the Person class, CORT improved HOTA (43.98 vs. 41.62) and MOTA (50.94 vs. 44.82), reducing ID switches (263 vs. 324) while maintaining a higher FPS (27.2 vs. 24.2). These results confirm that the proposed multi-class association and cascade matching strategies enhance both tracking robustness and computational efficiency.
In summary, these results highlight the robustness of the proposed solution in achieving precise object localization and reducing ID switches in challenging multi-class scenarios, without penalizing real-time performance.
4.4 Behavior under unstable detections
@@
This section examines the behavior of the standard and the multi-class approach in two critical situations in which the detection network could provide sporadic erroneous outputs in a frame sequence. To show the potential impact of such events on the performance of the tracker, consider a scenario similar to the one illustrated in Fig. 7, where the red car appears in the scene at time $t_{a}$, but it is misclassified as a bus at time $t_{a}+2$. Also assume that all the other objects are correctly classified across the sequence. For the sake of clarity, only the detection responsible for the problem is shown in all the frames.
@@
To better understand the negative impact of such a misclassification, Fig. 7 also reports the evolution of the Hungarian matrices for both the standard monolithic approach and the proposed class-based solution, highlighting for each frame how this specific detection affects the size of the matrices and, hence, the number of operations required to fill them. In particular, the colored rows represent the elements required in the cost matrix to include the track that must be created to match the missing detection, while the colored columns represent dummy values that need to be created in the cost matrix to meet the squareness constraint imposed by the Hungarian match. The red and blue colors are associated with the two different classes reported in the experiment. In particular, the red color is associated with the Bus class while the blue color is associated with the Car class.
At time $t_{a}$, there are five predictions and five detections assumed to be correctly tracked.
Under this assumption, at time $t_{a+2}$, when the car is misclassified as a bus, one track is missing a corresponding detection and there is a new detection that requires the creation of a new tentative track. To correctly create the entry for the new track, cost matrix must have dimension $6\times 6$ and, in such a context, the red row represents the space needed for creating the new track, while the blue column represents the space needed for creating a dummy detection for the car track at frame $t_{a}$, which now has no corresponding detection and is updated by using the Kalman filter prediction. The same rows and columns are reported also for CORT’s approach, which is clearly more resource efficient. Assuming an in_patience value of 5 frames, then at time $t_{a+5}$, when the car is again correctly recognized, the spatial resolution of the matrices does not change, meaning that the proposed approach not only benefits from the parallel computation, but also takes advantage of a better resources usage, reducing the number of operations to be executed within each thread. The transient between [$t_{a+2}$; $t_{a+5}$] shows that the just mentioned improvements not only apply for the frame where the misclassification occurs, but also for a time interval that depends on the tracking hyperparameters discussed in Section 3.4.
@@
A simple solution to attenuate the two problems described in this section is to increase the acceptance threshold of the detection network. Raising the value of such a threshold ensures that only reliable detections are passed to the tracker to compute the costs for the assignment problem, and it is very unlikely to have unstable detections with a high confidence score. At the same time, raising the threshold increases the number of false negatives. However, this is not a major issue, because the Kalman filter can efficiently compensate for some missing detections once the track is promoted effective, as explained in Section 3.4. This is the reason why in Table 2 the value for the precision (MOTP) is lower than that of the BYTEtrack, while the accuracy (MOTA) is higher because of the reduced number of id-switches.
5 Conclusion and Future work
@@
This work presented CORT, an efficient method to enhance the timing performance of modern real-time tracking algorithms. Such an enhancement is obtained by splitting the Hungarian matching algorithm into smaller association problems, carried out per class. Considering that the matching algorithm has an $O(n^{3})$ time complexity, where $n$ is the number of elements in the cost matrix, and that each sub-matrix can be solved in parallel on different cores, the proposed approach significantly improves the performance of state-of-the-art tracking-by-detection systems. In real-world applications, such as autonomous vehicles and surveillance, high-speed tracking of CORT enables a rapid and reliable identification of multiple objects belonging to different classes. This capability supports real-time decision-making in complex environments, where stable execution times and reduced resource demands are crucial. As discussed in Section 4.1, when the tracker operates in scenarios with a large number of objects, the proposed parallel solution demonstrates clear advantages in execution speed and scalability over monolithic approaches. This is also confirmed by the experimental results reported in Table 1, which show that the gap between the execution time of the two methods significantly increases in scenarios with a larger number of objects, underscoring the superior scalability of the proposed approach.
As a future work, we plan to further improve efficiecy by accelerating some tensor operations on FPGA, as those required by the Kalman filter and the Hungarian algorithm. Direct communication between modules on the FPGA will allow these tasks to be offloaded from the CPU, achieving more stable execution times and higher system throughput. Another optimization will be the dynamic tuning of the tracker’s hyperparameters. For example, re-identification thresholds could be adjusted in real-time based on scene complexity or object density, enabling the tracker to automatically adapt to changing conditions without manual intervention.
Acknowledgments
@@
This work was supported by the Department of Excellence of Robotics and AI, Scuola Superiore Sant’Anna, Pisa, Italy.